# Systematic study of confinement induced effects on atomic electronic structure

### v1

@@ 3 Computational Details
FEM offers a straightforward way to the numerical solution of the radial functions $R^{\sigma}_{nl}$;36 a review of the employed FEM approach can be found in refs. 37 and 38. In short, the radial domain is first divided into $N_{\mathrm{elem}}$ segments $r\in[r_{i}^{\mathrm{start}},r_{i}^{\mathrm{end}}]$ called elements. An exponential radial grid $\{r_{i}\}_{i=0}^{N_{\mathrm{elem}}}$ is employed, so that the size of the elements increases with the distance from the nucleus.37, 38 A piecewise polynomial basis of shape functions $B_{n}(r)$ is then built up in each element.37, 38 Finally, the numerical radial basis functions to be used in eq. 4 are built from FEM shape functions $B_{n}(r)$ as
@@
The endpoint of the last element is called the practical infinity, $r_{\infty}$, and all basis functions are built to vanish at $r_{\infty}$.37, 38 The physical interpretation of $r_{\infty}$ is that there is a hard-wall potential at this point; eq. 3 is therefore already built-in in the FEM approach, and this feature has been previously used in many studies, such as refs. 17 and 38. Thus, $r_{\infty}$ represents a physical parameter that expresses the location of the hard-wall boundary. In contrast, $r_{\infty}$ is a discretization parameter in studies of unconfined atoms (which represent the typical applications of FEM in the literature) that needs to be converged such that the obtained solution does not change if an even larger value is employed for $r_{\infty}$.
All calculations in this work were carried out with the free and open-source39 HelFEM program; the present implementation is publicly available on GitHub.40 Unless otherwise specified, the calculations employed shape functions defined by 15-node Lagrange interpolating polynomials (LIPs) specified by Gauss–Lobatto quadrature nodes; this corresponds to the use of a 14th order polynomial basis set. All calculations were converged to the complete basis set (CBS) limit with respect to the number of radial elements: radial elements were added until a $1\mu E_{h}$ precision in the total energy was achieved. We can thus be assured that we have found the numerically exact wave function.
We employ the optimal FEM implementation of the fractional occupation formalism recently described in refs. 28 and 29. All calculations employed the widely-used PBE exchange-correlation functional22, 23 as implemented in Libxc.41
We initiated the work by identifying the three lowest lying configurations of all elements with atomic number $1\leq Z\leq 54$ in the hard-wall potential for each $r_{c}\in\{1.0,1.2,1.4,\dots,10.0\}a_{0}$ with both spin-restricted and spin-polarized densities respectively, with a fixed angular momentum cutoff $l_{\mathrm{max}}=3$. The configurations were found with the automated algorithm introduced in ref. 28. The configurations considered for the spin-(un)restricted analysis were then chosen as the union of the low-lying configurations obtained from the spin-(un)restricted search over all of the above values of $r_{c}$.
@@ 4.1 Ionization energies of unconfined atoms
As a first step to validate our methods we compute ionization energies for the unconfined H–Xe atoms in the spin-restricted as well as the spin-polarized formalisms. The ionization energies can be calculated in two ways. First, the ionization energy is determined in the $\Delta$SCF method as the difference between the total energies of the neutral atom and its cation
$E_{+}=E_{\mathrm{atom}}-E_{\mathrm{ion}}.$ (7)
Second, while there is no Koopmans’ theorem42 for DFT as in Hartree–Fock, Janak’s theorem43 provides a justification for an analysis based on orbital energies.The ionization energy is then determined as
$E_{+}=\epsilon^{\mathrm{HOAO}},$ (8)
where $\epsilon^{\mathrm{HOAO}}$ is the eigenvalue of the highest occupied atomic orbital (HOAO).
(a) Ionization energies of the H–Xe atoms obtained with spin-restricted densities.
(b) Ionization energies of the H–Xe atoms obtained with spin-polarized densities.
Figure 1: Comparison of the ionization energies of gas-phase atoms computed with spin-restricted (fig. 1(a)) and spin-polarized (fig. 1(b)) densities in this work (TW) with $\Delta$SCF or Janak’s theorem against experimental values of Kramida et al.44.
The comparison of our computed values to the experimental values of Kramida et al.44 is depicted in fig. 1. The values obtained from Janak’s theorem are consistently and significantly lower than the experimental values and we will therefore not discuss them further. The $\Delta$SCF formalism produces ionization energies in good agreement with experiment. As expected, the results of the spin-polarized calculations are in better agreement with experiment than those from the spin-restricted calculations. Notable differences can be observed already for the H and He atoms. Spin-polarized calculations also faithfully reproduce the dip in the ionization energy beyond half filling of the $p$ shell, while spin-restricted calculations predict that the ionization energy increases monotonically with the filling of the $p$ shell.
@@ 4.2 Ionization radii
Having validated that the employed DFT method produces reliable ionization energies, we can proceed to computing ionization radii. The ionization radius $r_{+}$ is the value of $r_{c}$ for which it becomes energetically favorable for an atom to cast away its outermost electron. The analysis of Boeyens3 was based on non-relativistic Hartree–Fock calculations for $1\leq Z\leq 102$, and $r_{+}$ was identified via Koopmans’ theorem42 by finding the zero of the orbital energy of the outermost valence electron shell. However, as discussed in section 1, Boeyens3 did not employ a proper hard-wall potential, and did not consider the changing nature of the ground state configuration as a function of confinement.
@@ 4.2.1 $\Delta$SCF
The ionization radius $r_{+}$ for atom X is obtained within the $\Delta$SCF procedure by solving the nonlinear equation
$E({\rm X}^{+},r_{+})=E({\rm X},r_{+}),$ (9)
@@
[…]
The second method is to realize that confinement can affect the ground state configurations of the atom and and its cation, and to compute the energies $E(X,r_{c})$ and $E(X^{+},r_{c})$ for the lowest configurations at each $r_{c}$. The resulting ionization radii are shown in table 2.
H 1.95 1.86 He 1.35 1.40
Li 4.17 4.06 Be 3.14 3.21 B 3.06 3.01 C 2.67 2.58 N 2.38 2.28 O 2.15 2.27 F 1.97 2.02 Ne 1.82 1.83
Na 4.39 4.31 Mg 3.71 3.77 Al 4.14 4.09 Si 3.65 3.55 P 3.28 3.18 S 2.99 3.12 Cl 2.76 2.81 Ar 2.56 2.57
K 5.26 5.18 Ca 4.61 4.66 Sc 4.37 4.46 Ti 4.54 4.27 V 4.43 4.09 Cr 4.08 3.94 Mn 3.95 3.96 Fe 3.83 3.79 Co 3.73 3.65 Ni 3.63 3.58 Cu 3.54 3.51 Zn 3.31 3.34 Ga 4.07 4.02 Ge 3.72 3.63 As 3.45 3.34 Se 3.22 3.34 Br 3.03 3.08 Kr 2.86 2.87
Rb 5.57 5.50 Sr 4.98 5.03 Y 4.72 4.82 Zr 4.88 4.55 Nb 4.52 4.38 Mo 4.37 4.23 Tc 3.59 4.13 Ru 3.40 4.04 Rh 3.24 3.29 Pd 3.09 3.11 Ag 3.88 3.85 Cd 3.66 3.69 In 4.42 4.37 Sn 4.09 4.01 Sb 3.84 3.74 Te 3.62 3.74 I 3.44 3.49 Xe 3.28 3.29
Table 1: Ionization radii $R$ in $a_{0}$ for the H–Xe atoms from spin-restricted (in red) and spin-polarized (in blue) non-relativistic $\Delta$SCF calculations with fractional occupations with configurations fixed to the ground states of the unconfined neutral atom and its cation.
H 1.95 1.86 He 1.35 1.40
Li 4.17 4.06 Be 3.14 3.21 B 3.06 3.01 C 2.67 2.58 N 2.38 2.28 O 2.15 2.27 F 1.97 2.02 Ne 1.82 1.83
Na 4.39 4.31 Mg 3.71 3.77 Al 4.14 4.09 Si 3.65 3.55 P 3.28 3.18 S 2.99 3.12 Cl 2.76 2.81 Ar 2.56 2.57
K 5.26 5.18 Ca 4.82 4.78 Sc 4.62 4.49 Ti 4.41 4.14 V 4.14 3.82 Cr 3.89 3.94 Mn 3.68 3.83 Fe 3.50 3.72 Co 3.36 3.46 Ni 3.23 3.26 Cu 3.54 3.51 Zn 3.31 3.34 Ga 4.07 4.02 Ge 3.72 3.63 As 3.45 3.34 Se 3.22 3.34 Br 3.03 3.08 Kr 2.86 2.87
Rb 5.57 5.50 Sr 5.13 5.07 Y 4.86 4.59 Zr 4.41 4.16 Nb 4.07 3.84 Mo 3.81 4.17 Tc 3.59 3.79 Ru 3.40 3.51 Rh 3.24 3.29 Pd 3.09 3.11 Ag 3.88 3.85 Cd 3.66 3.69 In 4.42 4.37 Sn 4.09 4.01 Sb 3.84 3.74 Te 3.62 3.74 I 3.44 3.49 Xe 3.28 3.29
Table 2: Ionization radii $R$ in $a_{0}$ for the H–Xe atoms from spin-restricted (in red) and spin-polarized (in blue) non-relativistic $\Delta$SCF calculations with fractional occupations, considering the $R$ dependence of the ground state of the neutral atom and its cation.
The ionization radius for the hydrogen atom is an interesting place to start the analysis. After all, one would expect the largest error to be made for this atom, as its exact ground state is fully spin polarized, while our calculation is for an atom with $1/2$ spin-up and $1/2$ spin-down electrons. The corresponding exact model is also analytically solvable: Sommerfeld and Welker45 obtained the ionization radius $1.8352a_{0}$ in their pioneering study. It is quite reassuring that our spin-polarized DFT value of $1.86a_{0}$ deviates from this exact value for H by only 1.4%, while the spin-restricted value of $1.95a_{0}$ exhibits a much larger deviation of 6.3%—which is still remarkably accurate.
The comparison of the results of the first and second methods proves interesting. Although inclusion of the relaxation effects does not appear to be important for most of the studied elements, large differences are observed in the pre-$d$ and $d$ blocks of the periodic table (Ca–Ni and Sr–Mo).
[…]
However, significant changes in energy upon relaxing the configuration may also occur for the cation, in which case the ionization radius can increase. Interestingly, such increases when the ground states are relaxed are observed for the Ca, Sc and Sr atoms in both the spin-restricted and spin-polarized cases, as well as the Y atom in the spin-restricted case. The corresponding cations exhibit a ground state crossing that moves the cross-over point of eq. 9 down in energy, and thereby the ionization radius to the right; see fig. 2 for an illustration of these four exceptional cases.
@@
Figure 2: Ground state crossings for the spin-polarized Ca, Sc and Sr atoms, and the spin-restricted Y atom in the hard-wall potential. Configurations for the neutral atoms shown as solid lines, and those corresponding to the monocation as dashed lines.
The differences observed in ionization radii with and without relaxing the electronic configuration underline the importance of considering various possible configurations as a function of confinement. Many transition metal atoms and their monocations exhibit ground state crossings already at weak confinement, due to the low-lying $s$-$d$ excitations in these systems (see section 4.3).
The spin-polarized and spin-restricted ionization radii are different for all atoms, which is not surprising given that either the neutral atom or its cation is open-shell. Interestingly, the differences between these two approaches are the smallest for the noble gas atoms, while alkali metals show much larger differences between spin-polarized and spin-unpolarized calculations. As the spin-polarized calculations also gave ionization energies in good agreement with experiment in section 4.1, the spin-polarized ionization radii ¸should be the most accurate.
@@ 4.2.2 Janak’s theorem
Another possibility is to employ a procedure similar to Boeyens3 who used Koopmans’ theorem to determine the ionization radii. As discussed in section 4.1, we can use Janak’s theorem, and the ionization radius $r_{+}$ is now found from
$\epsilon^{\text{HOAO}}({\rm X},r_{+})=0$ (10)
where $\epsilon^{\text{HOAO}}({\rm X})$ is the energy of the HOAO of the neutral atom X, with the hard wall boundary at $r_{c}=r_{+}$.
We report ionization radii obtained from calculations using Janak’s theorem without the TSA as tables in the SI, as we already found above in section 4.1 that these calculations failed to reproduce ionization energies in a reliable manner.
We note in passing that an exact solution to eq. 10 is problematic in some cases that exhibit a ground state crossing in the relevant range for $r_{+}$. A case in point is the V atom: the value of $r_{+}$ corresponds to the location of a ground state crossing between a state with $\epsilon^{\mathrm{HOAO}}(r)<0$ and a state with $\epsilon^{\mathrm{HOAO}}(r)>0$, which means that eq. 10 is not truly satisfied no matter how tightly $r_{+}$ is converged.
The origin of this problem is the present use of electron configurations with integer occupations. However, it has long been known that the use of non-integral occupation numbers for the $4s$ and $3d$ shells can result in a lower total energy for transition metal atoms in average-of-configurations Hartree–Fock calculations.46 This phenomenon is also topical for the present calculations with the PBE functional,47, 28 and has also been found in DFT calculations that do not employ any exchange-correlation functional.48 Calculations employing variationally optimized fractional occupation numbers would allow finding an exact solution to eq. 10. They are also expected to lead to somewhat different solutions of eq. 9. However, we do not consider such fractional occupations in this work, since integer occupation numbers are more commonly used and allow a clearer physical interpretation of the results.
Sen et al.13 computed atomic ionization radii with Janak’s theorem for the He–Ne atoms with exchange-only calculations based on Becke’s 1988 generalized-gradient functional.49 In contrast to our work, Sen et al.13 employed the transition state approximation (TSA) of Slater et al.46 in their calculations; the orbital energies were thus determined in calculations where half an electron was removed from the highest-lying orbital. Sen et al.13 compared this method to $\Delta$SCF calculations and found that the two methods yield ionization radii in close agreement to each other. Despite the differences between our approaches, we will see below that our results are in good agreement with those of Sen et al.13.

### v2

@@ 3 Computational Details
FEM offers a straightforward way to the numerical solution of the radial functions $R^{\sigma}_{nl}$;52 a review of the employed FEM approach can be found in refs. 53 and 54. In short, the radial domain is first divided into $N_{\mathrm{elem}}$ segments $r\in[r_{i}^{\mathrm{start}},r_{i}^{\mathrm{end}}]$ called elements. An exponential radial grid $\{r_{i}\}_{i=0}^{N_{\mathrm{elem}}}$ is employed, so that the size of the elements increases with the distance from the nucleus.53, 54 A piecewise polynomial basis of shape functions $B_{n}(r)$ is then built up in each element.53, 54 Finally, the numerical radial basis functions to be used in eq. 4 are built from FEM shape functions $B_{n}(r)$ as
@@
The endpoint of the last element is called the practical infinity, $r_{\infty}$, and all basis functions are built to vanish at $r_{\infty}$.53, 54 The physical interpretation of $r_{\infty}$ is that there is a hard-wall potential at this point; eq. 3 is therefore already built-in in the FEM approach, and this feature has been previously used in many studies, such as refs. 19 and 54. Thus, $r_{\infty}$ represents a physical parameter that expresses the location of the hard-wall boundary. In contrast, in studies of unconfined atoms (which represent the typical applications of FEM in the literature), $r_{\infty}$ is a discretization parameter that needs to be converged such that the obtained solution does not change if an even larger value is employed for $r_{\infty}$.
As our main focus is on atoms in confinement, the calculations unconfined atoms employed the default value $r_{\infty}=40a_{0}$, which is sufficient to reproducing the complete basis set limit energy of the ground states of neutral atoms and their cations. We note here that as larger values of $r_{\infty}$ may be necessary to capture the behavior of loosely bound states, energies of the excited states in unconfined atoms may not be converged to the complete basis set limit with respect to this parameter. In contrast, the energies for the excited states in the confined atoms are converged to the complete basis set limit, as is explained below.
All calculations in this work were carried out with the free and open-source55 HelFEM program; the present implementation is publicly available on GitHub.56 The calculations employed shape functions defined by 15-node Lagrange interpolating polynomials (LIPs) specified by Gauss–Lobatto quadrature nodes; this corresponds to the use of a 14th order polynomial basis set. All calculations were converged to the complete basis set (CBS) limit with respect to the number of radial elements: radial elements were added until a $1\mu E_{h}$ precision in the total energy was achieved. We can thus be assured that we have found the numerically exact wave function.
We employ the optimal FEM implementation of the fractional occupation formalism recently described in refs. 37 and 38. We perform calculations within the local density approximation57, 58 employing the Perdew–Wang (PW92) correlation functional,59 within the generalized-gradient approximation (GGA) employing the Perdew–Burke–Ernzerhof (PBE) exchange-correlation functional,32, 33 and within the meta-GGA approximation employing the r2SCAN exchange-correlation functional,60, 61 all as implemented in Libxc62 using the lda_x-lda_c_pw, gga_x_pbe-gga_c_pbe, and mgga_x_r2scan-mgga_c_r2scan keywords, respectively.
We initiated the work by identifying the three lowest lying configurations of all elements with atomic number $1\leq Z\leq 54$ in the hard-wall potential for each $r_{c}\in\{1.0,1.2,1.4,\dots,10.0\}a_{0}$ with both spin-restricted and spin-polarized densities, respectively, with a fixed angular momentum cutoff $l_{\mathrm{max}}=3$. The configurations were determined with the automated algorithm introduced in ref. 37. The configurations considered for the spin-(un)restricted analysis were then chosen as the union of the low-lying configurations obtained from the spin-(un)restricted search over all of the above values of $r_{c}$.
@@
The configuration search above was performed with the PBE functional, only. However, as we shall see in section 4, the results are qualitatively independent of the employed density functional approximation. For this reason, we are confident that our search yielded all relevant configurations for the PW92 and r2SCAN functionals, as well.
@@ 4.1 Ionization energies of unconfined atoms
As a first step to validate our methods we compute ionization energies for the unconfined H–Xe atoms in the spin-restricted and spin-polarized formalisms according to the two methods discussed in section 2.2.
(a) Ionization energies of the H–Xe atoms obtained with the PBE functional and spin-restricted densities.
(b) Ionization energies of the H–Xe atoms obtained with the PBE functional and spin-polarized densities.
Figure 1: Comparison of the ionization energies of unconfined atoms computed with spin-restricted (fig. 1(a)) and spin-polarized (fig. 1(b)) densities in this work (TW) with $\Delta$SCF or Janak’s theorem against experimental values of Kramida et al.63.
As the three dissimilar functionals yield ionization energies that are all in good agreement with the experimental data, we only discuss the PBE data here. The comparison of our PBE values to the experimental values of Kramida et al.63 is depicted in fig. 1; analogous plots for the PW92 and r2SCAN functionals are included in the SI.
The values obtained from Janak’s theorem are consistently and significantly lower than the experimental values and we will therefore not discuss them further.
The $\Delta$SCF formalism produces ionization energies in good agreement with experiment. As expected, the results of the spin-polarized calculations are in better agreement with experiment than those from the spin-restricted calculations. Notable differences can be observed already for the H and He atoms. Spin-polarized calculations also faithfully reproduce the dip in the ionization energy beyond half filling of the $p$ shell, while spin-restricted calculations predict that the ionization energy increases monotonically with the filling of the $p$ shell.
@@ 4.2.1 $\Delta$SCF
H 1.95 1.85 Ca 4.60 4.66 Sc 4.37 4.46 Ti 4.53 4.27 V 4.43 4.09 Cr 4.08 3.94 Mn 3.94 3.95 Fe 3.83 3.79 Co 3.72 3.65 Ni 3.63 3.57 He 1.34 1.40
Li 4.17 4.06 Be 3.14 3.21 Sr 4.97 5.02 Y 4.72 4.82 Zr 4.87 4.55 Nb 4.52 4.37 Mo 4.37 4.22 Tc 3.58 4.12 Ru 3.39 4.04 B 3.05 3.01 C 2.66 2.58 N 2.37 2.28 O 2.15 2.27 F 1.96 2.01 Ne 1.81 1.83
Na 4.38 4.30 Mg 3.70 3.76 Al 4.13 4.08 Si 3.64 3.55 P 3.28 3.18 S 2.99 3.11 Cl 2.75 2.80 Ar 2.56 2.57
K 5.26 5.18 Ca 4.82 4.77 Sc 4.62 4.49 Ti 4.40 4.14 V 4.13 3.82 Cr 3.88 3.94 Mn 3.67 3.83 Fe 3.50 3.72 Co 3.35 3.46 Ni 3.22 3.25 Cu 3.54 3.50 Zn 3.30 3.33 Ga 4.06 4.01 Ge 3.71 3.62 As 3.44 3.34 Se 3.21 3.33 Br 3.02 3.07 Kr 2.86 2.87
Rb 5.56 5.49 Sr 5.12 5.07 Y 4.85 4.59 Zr 4.41 4.15 Nb 4.07 3.83 Mo 3.80 4.17 Tc 3.58 3.79 Ru 3.39 3.51 Rh 3.23 3.29 Pd 3.09 3.10 Ag 3.88 3.84 Cd 3.66 3.69 In 4.41 4.37 Sn 4.09 4.00 Sb 3.83 3.73 Te 3.62 3.73 I 3.43 3.48 Xe 3.27 3.28
Table 1: Ionization radii $R$ in $a_{0}$ for the H–Xe atoms from spin-restricted (in red) and spin-polarized (in blue) non-relativistic $\Delta$SCF calculations with fractional occupations and the PBE functional with configurations fixed to the ground states of the unconfined neutral atom and its cation. Values obtained while considering the $R$ dependence of the ground state of the neutral atom and its cation are highlighted in the insert for the elements where the value differs from the static configurations.
As discussed in section 2.3, $\Delta$SCF ionization radii can be computed in two ways. In the first method we fix the configurations to those of the ground state of the neutral atoms and their cations, respectively. In the second method we relax the configurations of the atoms and cations to the lowest configurations at each radius. The resulting ionization radii for the H–Xe atoms obtained with the PBE functional are shown in table 1 and its insert, respectively, when the second method predicts a different ionization radius than the fixed configuration approach.
The ionization radius for the hydrogen atom is an interesting place to start the analysis. The corresponding exact model is analytically solvable: Sommerfeld and Welker64 obtained the ionization radius $1.8352a_{0}$ in their pioneering study. The exact ground state of hydrogen is fully spin polarized and our spin-polarized PW92, PBE and r2SCAN values of $1.90a_{0}$, $1.86a_{0}$, and $1.84a_{0}$ deviate from the exact value by only 4.5%, 1.4% and 0.3%, respectively. The spin-restricted values of $1.99a_{0}$, $1.95a_{0}$, and $1.96a_{0}$ exhibit much larger deviations of 8.4%, 6.3%, and 6.8 %, respectively, but are still remarkably accurate considering that our spin-restricted calculations are for an atom with $1/2$ spin-up and $1/2$ spin-down electrons.
Moving on to heavier elements, the comparison of the results from the fixed electron configuration and relaxed configuration methods proves interesting. Although inclusion of the relaxation effects does not appear to be important for most of the studied elements, large differences are observed in the pre-$d$ and $d$ blocks of the periodic table (Ca–Ni and Sr–Ru).
@@
However, significant changes in energy upon relaxing the configuration may also occur for the cation, in which case the ionization radius can increase. Interestingly, such increases when the ground states are relaxed are observed for the Ca, Sc, and Sr atoms in both the spin-restricted and spin-polarized cases, as well as the Y atom in the spin-restricted case. The corresponding cations exhibit a ground state crossing that moves the cross-over point of eq. 8 down in energy, and thereby the ionization radius to the right; see fig. 2 for an illustration of these four exceptional cases.
Because the various configurations’ energies span several orders of magnitude under the studied range of confinement radii, we compare the energy of the confined atom with electron configuration $i$ to that of the ground state configuration of the unconfined atom
$\Delta E_{i}(r_{c})=E_{i}(r_{c})-E_{0}^{\mathrm{unconfined}},$ (14)
where $i$ is the state in question in all plots in this study.
@@
Figure 2: Ground state crossings for the spin-polarized Ca, Sc and Sr atoms, and the spin-restricted Y atom in the hard-wall potential with the PBE functional. Configurations for the neutral atoms shown as solid lines, and those corresponding to the monocation as dashed lines.
The differences observed in ionization radii with and without relaxing the electron configuration underline the importance of considering all possible low lying configurations as a function of confinement. Many transition metal atoms and their monocations exhibit ground state crossings already in weak confinement, due to the low-lying $s$-$d$ excitations in these systems (see section 4.4).
The spin-polarized and spin-restricted ionization radii differ for all atoms, which is not surprising given that either the neutral atom or its cation is going to be open-shell. Interestingly, the differences between these two approaches are the smallest for the noble gas atoms, while alkali metals show much larger differences between spin-polarized and spin-unpolarized calculations. As the spin-polarized calculations gave ionization energies in best agreement with experiment in section 4.1, the spin-polarized ionization radii should be the most accurate.
Figure 3: Ionization radii for the H–Xe atoms with the PW92, PBE and r2SCAN functionals obtained from spin-polarized densities via $\Delta$SCF with relaxed configurations.
A comparison of the ionization radii obtained with the PW92, PBE and r2SCAN functionals with spin-polarized densities and relaxed configurations is shown in fig. 3. Except for the alkali metals, as well as the Cr, Y, and Nb atoms, the dissimilar density functional approximations are in excellent agreement. We further validate this agreement by qualitatively studying the functional dependence of the behavior of all low lying configurations of the Ca atom in figs. 5, 4 and 6, while the graphical representation of the results of the other atoms is included in the SI. We chose to feature Ca here due to its interesting chemistry which will be discussed below in section 4.4. The only major difference between the functionals is that r2SCAN predicts the $4s^{1}3d^{1}$ configuration to be lower in energy than the $4s^{1}4p^{1}$ configuration in the unconfined atom, while PW92 and PBE predict the opposite.
Despite such minor differences in the ordering of some states, or relative energy differences between states, the qualitative behaviors obtained with the three different density functional approximations agree well with each other; the changes in the electron configurations are practically independent of the functional. This suggests that the phenomena found with the methodology employed in this work are not artefacts of the used level of theory, but true physical phenomena within the studied model of confinement.
Figure 4: Energies of low lying configurations of hard-wall confined spin-polarized Ca computed with PW92 shown as the energy difference from unconfined Ca as a function of the confinement radius. Note semilogarithmic scale.
Figure 5: Energies of low lying configurations of hard-wall confined spin-polarized Ca computed with PBE shown as the energy difference from unconfined Ca as a function of the confinement radius. Note semilogarithmic scale.
Figure 6: Energies of low lying configurations of hard-wall confined spin-polarized Ca computed with r2SCAN shown as the energy difference from unconfined Ca as a function of the confinement radius. Note semilogarithmic scale.
@@ 4.2.2 Janak’s theorem
We report ionization radii obtained from calculations using Janak’s theorem as tables in the SI, as we already found above in section 4.1 that these calculations failed to reproduce ionization energies in a reliable manner.
Sen et al.15 computed atomic ionization radii with Janak’s theorem for the He–Ne atoms with exchange-only calculations based on Becke’s 1988 generalized-gradient functional.65 In contrast to our work, Sen et al.15 employed the transition state approximation (TSA) of Slater et al.47 in their calculations; the orbital energies were thus determined in calculations where half an electron was removed from the highest-lying orbital. Sen et al.15 compared this method to $\Delta$SCF calculations and found that the two methods yield ionization radii in close agreement to each other. Despite the differences between our approaches, we will see below that our results are in good agreement with those of Sen et al.15.